OpenStudy (anonymous):

5/6 divide by 1/2* 1/4 (order of operation).... can you show the steps?

OpenStudy (anonymous):

= (5/6)/(1/2[1/4]) = (5/6)/(1/8) = 5/6(8) = 40/6 or (40/2)/(6/2) = 20/3 or 6 2/3